1 Evaluate the integral 6 -sin ل dt. 6 T6 -sin dt = | 한
Answer #1

sel SA sin -3 3 ) dt sin (4-3) de Let. -3=4 ts - sinuda ²-[-cosu] tc. -3) ť6-3=u -6t7dt=du -6 dt = du 47 COS te Ans 76If you have any doubt regarding this please let me know. If you understand the solution then please give me a thumbs up. Thanks!!